OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form. Choose one answer. a. 3x - 3y = -8 b. x - 3y = -9 c. x - 3y = -8 d. 3x + y = 9

OpenStudy (aroub):

ok do u know how to find the slope?

OpenStudy (saifoo.khan):

m = (4-3)/(4-1) m = 1/3

OpenStudy (saifoo.khan):

4 = 4/3 + c c = 8/3

OpenStudy (saifoo.khan):

y = mx + c 3y = x + 8

OpenStudy (saifoo.khan):

your answer: c. x - 3y = -8

OpenStudy (aroub):

ty saif! then u choose any of the points u have ill choose (4,4) y=mx+b y=8/3 the answer is c. x - 3y = -8